# Bayesian Optimization with Preference Exploration using a Monotonic Neural Network Ensemble

### v2

@@ 4 Utility Surrogate Model and Pairwise Comparison Acquisition Function
The problem described above was originally proposed by Lin et al. (2022), and they handle it in a BO framework. This paper will follow their framework (albeit with different assumptions about permissible outputs for the Preference Exploration stage) but adjust the surrogate model and the corresponding acquisition function.
@@
In the Preference Exploration stage, the output pair is chosen according to an acquisition function called the Expected Utility of the Best Option (EUBO) (Lin et al., 2022). EUBO is defined as $\operatorname{EUBO}\left(y_{1,t},y_{2,t}\right)=\mathbb{E}_{m}\left[\max\left\{g\left(y_{1,t}\right),g\left(y_{2,t}\right)\right\}\right],$ where the expectation is over the posterior of the utility surrogate model $g$ at the time the query is chosen. Lin et al. (2022) assume that the experimenter can show the DM any pair for their preference decision even if the solutions do not exist. In this case, it is not straightforward to determine the pair using EUBO because in the acquisition optimization, the search range of the output $y_{1,t}$, $y_{2,t}$ is unknown. Hence, to find the pair of outputs, EUBO needs to be transformed as $\operatorname{EUBO}\left(x_{1,t},x_{2,t}\right)=\mathbb{E}_{m}\left[\max\left\{g\left(f(x_{1,t})\right),g\left(f(x_{2,t})\right)\right\}\right]$ and find the pair $(x_{1,t},x_{2,t})\in\mathcal{X}^{2}$, then use a posterior sample of $(f(x_{1,t}),f(x_{2,t}))$ as the comparison pair $(y_{1,t},y_{2,t})$.
@@ 4.1 Monotonic Neural Network Ensemble as Utility Surrogate Model
We propose using a neural network as the surrogate model. The advantages of neural networks are manifold. First, neural networks with sufficient capacity can approximate complex utility functions without requiring explicit assumptions about their functional form, providing the desired modeling flexibility. Secondly, by transforming the weights into positive space, we ensure the model is monotonically increasing, as suggested by Sill (1997). Specifically, we will apply an exponential transformation to the weights, guaranteeing that all weights are positive. The following theorem states that a monotonic neural network is able to model any bounded monotonically increasing function with bounded partial derivatives:
@@
(Sill (1997)) Let $m(x)$ be any continuous, bounded monotonic function with bounded partial derivatives, mapping $0,1]^{d}$ to $\mathcal{R}$. Then there exists a function $m_{net}(x)$ which can be implemented by a monotonic network and is such that, for any $\epsilon$ and any $x\in[0,1]^{d},\left|m(x)-m_{net}(x)\right|<\epsilon$.
@@
where $p_{i}=\pm 1$ is the corresponding preference of $y_{1,i}$ vs. $y_{2,i}$, and $\alpha$ is a positive parameter that is learned during training. Additionally, while the BOPE problem setting primarily deals with binary preferences ($\pm 1$), we can extend it to handle tied preferences (denoted by $p_{i}=0$). The Hinge Loss can be modified accordingly: $\mathcal{L}(\mathcal{P}_{t})=\sum_{(y_{1,i},y_{2,i})\in\mathcal{P}t}[1_{p_{i}\neq 0}\cdot\max\{0,1-\alpha\cdot(g(y_{1,i})-g(y_{2,i}))\cdot p_{i}\}+1_{p_{i}=0}\cdot|g(y_{1,i})-g(y_{2,i})|]$. Lastly, we employ neural network ensembles (Hansen and Salamon, 1990; Lakshminarayanan et al., 2017) to obtain uncertainty estimates that are important for the acquisition criterion.
@@
To incorporate uncertainty quantification, we explored several established methods, primarily focusing on Bayesian Neural Networks (BNNs) and ensemble techniques. While BNNs trained by Hamiltonian Monte Carlo (HMC) are generally superior for standard supervised learning with direct input-output observations $(X,Y)$ and performs well in ‘vanilla’ BO (see Figure 5 in Appendix A.2), it encounters significant challenges in preference learning. Briefly, the independent prior distribution of parameters hinders training due to scale invariance in preference learning (because the training set only shows relative comparisons of function values rather than absolute values). We refer to Appendix A.3 for more details. Also, given that Monte Carlo Dropout underperforms compared to ensemble methods (Lakshminarayanan et al., 2017), we focus the investigation on ensemble approaches and BNNs trained with variational inference (VI), particularly recommending ensemble methods for uncertainty quantification in preference learning. This recommendation is supported by several observations: recent studies have shown that ensembles provide more reliable uncertainty estimates than BNNs (Lakshminarayanan et al., 2017; Ovadia et al., 2019). In the specific context of Bayesian Optimization for Neural Architecture Search, White et al. (2021) found ensemble methods to outperform Bayesian estimation approaches, and generally the ensemble also shows a comparable but marginally lower performance than HMC-trained BNNs in ‘vanilla’ BO (Li et al., 2024). Finally, ensembles offer practical advantages through their inherent parallelizability, substantially reducing computational overhead, and as we will demonstrate later, naturally can be combined with normalization to handle scale invariance in preference learning.
@@
In a neural network ensemble, $M_{e}$ ($M_{e}>1$) base neural networks, denoted as $g^{j}~(j=1,...,M_{e})$, are simultaneously trained with the same training set but from different initial parameters. The predicted mean and variance are derived by aggregating outputs from these $M_{e}$ neural networks. It should be noticed that the output scales of the $M_{e}$ models might be different due to scale invariance. To handle this issue, we can normalize neural networks to ensure uniform scaling. Specifically, the output $g^{j}(x)$ is scaled as $\frac{g^{j}(x)-g^{j}_{\text{max}}}{\sigma^{j}}$, where $g^{j}_{\text{max}}$ represents the largest estimated utility value of the compared outputs: $g^{j}_{\text{max}}=max_{y\in\{y_{i}\mid(y_{1,i},y_{2,i})\in\mathcal{P}_{m},y_{i}\in\{y_{1,i},y_{2,i}\}}\{g^{j}(y)\}$ and $\sigma^{j}$ denotes the normalized deviation of compared outputs: $\sigma^{j}=\sqrt{\frac{\sum_{y\in\{y_{i}\mid(y_{1,i},y_{2,i})\in\mathcal{P}_{m},y_{i}\in\{y_{1,i},y_{2,i}\}\}}(g^{j}(y)-\mu^{j})^{2}}{|\mathcal{P}_{m}|}}$.
@@
Figure 2 shows two surrogate models trained with the same data: a GP and our MoNNE (detailed descriptions of these models are provided in Section 5). The left-hand side of Figure 2 displays two distinct 2D utility functions. The first utility (top) is a linear utility function defined as $g_{true}((y_{1},y_{2}))=y_{1}+1.5y_{2}$ and the second (bottom) is more complicated, defined as $g_{true}((y_{1},y_{2}))=\sqrt{y_{1}}+0.9\text{sin}(y_{1})+\ln(y_{2}+e^{y_{1}})+4.5\sqrt{y_{2}}$. Both functions increase monotonically in the range $0,10]^{2}$. For clearer visualization, we present slices of these functions. The $x$-axis represents either $y_{1}$ (value of the first dimension) or $y_{2}$ (value of the second dimension), as indicated in the legend. For the $y_{1}+y_{2}=10$’ slice, the $x$-axis shows $y_{1}$ values while $y_{2}=10-y_{1}$. For $y_{2}=2.5$’, the $x$-axis represents $y_{1}$ values with $y_{2}$ fixed at $2.5$. Similarly, for $y_{1}=2.5$’, the $x$-axis shows $y_{2}$ values with $y_{1}$ fixed at $2.5$. Using 15 randomly chosen pairs from $y_{1}+y_{2}=10$, we constructed the GP and MoNNE models. The middle plot illustrates the GP surrogate model, while the right plot depicts the MoNNE surrogate model. As evident in Figure 2, the MoNNE model more accurately captures the underlying utility function with only 15 output pairs. Furthermore, by incorporating monotonicity constraints, MoNNE maintains a strictly increasing relationship for legends $y_{1}=2.5$’ and $y_{2}=2.5$’, whereas the GP model exhibits non-monotonic behavior.
@@ 4.2.1 Experimentation
After building the two surrogate models $f$ and $g$, we adopt qNEIUU (Astudillo and Frazier, 2020; Lin et al., 2022) to choose the next solution to be evaluated for the Experimentation stage. Following Lin et al. (2022), we calculate the qNEIUU by Monte Carlo simulation
@@ 4.2.2 Preference Exploration
Our experiments demonstrate that IEUBO generally outperforms EUBO in the presence of utility noise. Detailed experimental results are provided in Appendix A.8. This performance difference may be attributed to how noise affects decision-maker preferences: under noisy conditions, the base model of MoNNE might learn incorrect correlations between outputs, potentially making the independence assumption of IEUBO more robust than EUBO’s consideration of correlations. Hence, we use IEUBO as the acquisition function.
@@ 5 Experimental Results
In this section, we compare the proposed method with benchmark algorithms on five test problems. The performance of different algorithms is measured by simple regret, defined as the gap between the true optimal utility and the best utility over the observations: $\text{max}_{x\in\mathcal{X}}\{g_{\text{true}}\left(f_{\text{true}}(x)\right)\}-\text{max}_{x\in{\{x_{i}|(x_{i},y_{i})\in D_{n}\}}}\{g_{\text{true}}\left(f_{\text{true}}(x)\right)\}$.
We use all four multi-objective problems that were used as benchmarks in (Lin et al., 2022), plus two additional ones. We test a wide range of utility functions from Lin et al. (2022); Astudillo and Frazier (2020), plus the Cobb-Douglas function, which is a standard function in economics, and LinearExponential, a sum of a linear component and an exponential product. Details can be found in Appendix A.5. Following the experiment setting of Lin et al. (2022), the initial number of output observations, that is, the initial dataset, is 16 (for $d<5$) or 32 (for $d\geq 5$), and the number of initial pairwise comparisons is 10 (for $d<5$) or 20 (for $d\geq 5$). In each iteration, we evaluate one solution, observing $f_{\text{true}}(x)$ and ask the DM for their preference over one pair of outputs. We compare the proposed method against the following benchmark algorithms: Random, qNParEGO (Daulton et al., 2020), qNEHVI (Daulton et al., 2020), PBO Astudillo et al. (2023), BOPE-Linear (using a linear model as the utility surrogate model), BOPE-GP (Lin et al., 2022), BOPE-MoNNE (Ours), BOPE-BMNN (Ours, using Variational Inference-trained Bayesian Monotonic Neural Network), Known Utility (the utility function is known, so the problem becomes a Composite BO (Astudillo and Frazier, 2019)). For more details about benchmark algorithms, see Appendix A.5.
We allow for noise in DM’s decision, as discussed in Section 3. As default, $\sigma_{\text{noise}}$ is set to $0.1$ (noise is a setting of the test problem, not the algorithm). Although $0.1$ may seem small, considering that BOPE-MoNNE can achieve regret less than $0.001$ as shown below in Figure 3, $0.1$ utility noise can cause enough trouble (Figure 21 in Appendix A.9 shows how many preference errors $0.1$ utility noise can make). We also conduct tests with $\sigma_{\text{noise}}=0$ and $\sigma_{\text{noise}}=0.5$ to evaluate how this noise affects the performance of MoNNEs and GP surrogate models, and the result is shown in Figure 22 in Appendix A.9.
@@
Figure 3 compares the performance of different methods, with subplots titled ‘A+B’ (A: output function, B: utility function). Excluding Known Utility algorithm, among all the BOPE algorithms, BOPE-MoNNE consistently achieves the best performance, followed by BOPE-BMNN and BOPE-GP, aligning with the uncertainty quality of these models reported in Table 1. We evaluated uncertainty quality using a modified cross-entropy loss function: $\mathbbm{1}_{p=1}\cdot\mathbb{P}(p=1)+\mathbbm{1}_{p=-1}\cdot\mathbb{P}(p=-1)$, where $\mathbb{P}(p=1)$ and $\mathbb{P}(p=-1)$ represent the model’s predicted probabilities for the DM’s preferences. This metric takes into account both the accuracy of the model and its uncertainty, and the larger the unit, the better. From Table 1, MoNNE exhibits the highest uncertainty quality, with BNN following and GP ranking last. It should be noticed that the poor uncertainty quality of GPs is mainly due to scale invariance, while the GP prediction accuracy is still good, although lower than MoNNE and BNN (see Table 2). The experiment details about these tables are reported in Appendix A.5. In addition, BOPE-Linear works slightly better than BOPE-MONNE when the utility function is linear, which is not surprising. However, in the other problems, the MoNNE model consistently outperforms the linear model.
Furthermore, we find that constraining output comparisons to observed values diminishes BOPE-GP’s performance (see Figure 19 in Appendix A.8). While removing this constraint improves BOPE-GP’s effectiveness, it still underperforms compared to BOPE-MoNNE.

### v3

@@ 4 Utility Surrogate Model and Pairwise Comparison Acquisition Function
The problem described above was originally proposed by Lin et al. (2022), and they handle it in a BO framework. This paper follows their framework (albeit with different assumptions about permissible outputs for the Preference Exploration stage) but adjusts the surrogate model and the corresponding acquisition function.
@@
In the Preference Exploration stage, the output pair is chosen according to an acquisition function called the Expected Utility of the Best Option (EUBO) (Lin et al., 2022). EUBO is defined as $\operatorname{EUBO}\left(y_{1,t},y_{2,t}\right)=\mathbb{E}_{m}\left[\max\left\{g\left(y_{1,t}\right),g\left(y_{2,t}\right)\right\}\right],$ where the expectation is over the posterior of the utility surrogate model $g$ at the time the query is chosen. Lin et al. (2022) assume that the experimenter can show the DM any pair for their preference decision even if the solutions do not exist. In this case, it is not straightforward to determine the pair using EUBO because in the acquisition optimization, the search range of the output $y_{1,t}$, $y_{2,t}$ is unknown. Hence, to find the pair of outputs, EUBO needs to be transformed as $\operatorname{EUBO}\left(x_{1,t},x_{2,t}\right)=\mathbb{E}_{m}\left[\max\left\{g\left(f(x_{1,t})\right),g\left(f(x_{2,t})\right)\right\}\right]$ to find the pair $(x_{1,t},x_{2,t})\in\mathcal{X}^{2}$, and the experimenter can use a posterior sample of $(f(x_{1,t}),f(x_{2,t}))$ as the comparison pair $(y_{1,t},y_{2,t})$.
@@ 4.1 Monotonic Neural Network Ensemble as Utility Surrogate Model
We propose using a neural network as the surrogate model. The advantages of neural networks are manifold. First, neural networks with sufficient capacity can approximate complex utility functions without requiring explicit assumptions about their functional form, providing the desired modelling flexibility. Second, by transforming the weights into positive space, we ensure the model is monotonically increasing, as suggested by Sill (1997). Specifically, we will apply an exponential transformation to the weights, guaranteeing that all weights are positive. The following theorem states that a monotonic neural network is able to model any bounded monotonically increasing function with bounded partial derivatives:
@@
(Sill (1997)) Let $m(x)$ be any continuous, bounded monotonic function with bounded partial derivatives, mapping $0,1]^{d}$ to $\mathbb{R}$. Then there exists a function $m_{net}(x)$ which can be implemented by a monotonic network and is such that, for any $\epsilon$ and any $x\in[0,1]^{d},\left|m(x)-m_{net}(x)\right|<\epsilon$.
@@
where $p_{i}=\pm 1$ is the corresponding preference of $y_{1,i}$ vs. $y_{2,i}$, and $\alpha$ is a positive parameter that is learned during training. Additionally, while the BOPE problem setting primarily deals with binary preferences ($\pm 1$), we can extend it to handle tied preferences (denoted by $p_{i}=0$). The Hinge loss can be modified accordingly: $\mathcal{L}(\mathcal{P}_{t})=\sum_{(y_{1,i},y_{2,i})\in\mathcal{P}_{t}}[\mathbbm{1}_{p_{i}\neq 0}\cdot\max\{0,1-\alpha\cdot(g(y_{1,i})-g(y_{2,i}))\cdot p_{i}\}+\mathbbm{1}_{p_{i}=0}\cdot|g(y_{1,i})-g(y_{2,i})|]$. Lastly, we employ neural network ensembles (Hansen and Salamon, 1990; Lakshminarayanan et al., 2017) to obtain uncertainty estimates that are important for the acquisition criterion.
@@
To incorporate uncertainty quantification, we explored several established methods, primarily focusing on Bayesian Neural Networks (BNNs) and ensemble techniques. While BNNs trained by Hamiltonian Monte Carlo (HMC) are generally superior for standard supervised learning with direct input-output observations $(X,Y)$ and performs well in ‘vanilla’ BO (see Figure 5 in Appendix A.2), it encounters significant challenges in preference learning. Briefly, the independent prior distribution of parameters hinders training due to scale invariance in preference learning (because the training set only shows relative comparisons of function values rather than absolute values). We refer to Appendix A.3 for more details. Also, given that Monte Carlo Dropout underperforms compared to ensemble methods (Lakshminarayanan et al., 2017), we focus the investigation on ensemble approaches and BNNs trained with variational inference (VI), particularly recommending ensemble methods for uncertainty quantification in preference learning. This recommendation is supported by several observations: recent studies have shown that ensembles provide more reliable uncertainty estimates than BNNs (Lakshminarayanan et al., 2017; Ovadia et al., 2019). In the specific context of BO for Neural Architecture Search, White et al. (2021) found ensemble methods to outperform Bayesian estimation approaches, and generally ensembles also show a comparable but marginally lower performance than HMC-trained BNNs in ‘vanilla’ BO (Li et al., 2024). Finally, ensembles offer practical advantages through their inherent parallelizability, substantially reducing computational overhead, and as we will demonstrate later, naturally can be combined with normalization to handle scale invariance in preference learning.
@@
In a neural network ensemble, $M_{e}$ ($M_{e}>1$) base neural networks, denoted as $g^{j}~(j=1,...,M_{e})$, are simultaneously trained with the same training set but from different initial parameters. The predicted mean and variance are derived by aggregating outputs from these $M_{e}$ neural networks. It should be noticed that the output scales of the $M_{e}$ models might be different due to scale invariance. To handle this issue, we can normalize neural networks to ensure uniform scaling. Specifically, the output $g^{j}(x)$ is scaled as $\frac{g^{j}(x)-g^{j}_{\text{max}}}{\sigma^{j}}$, where $g^{j}_{\text{max}}$ represents the largest estimated utility value of the compared outputs: $g^{j}_{\text{max}}=\text{max}_{y\in\{y_{i}\mid(y_{1,i},y_{2,i})\in\mathcal{P}_{m},y_{i}\in\{y_{1,i},y_{2,i}\}}\{g^{j}(y)\}$ and $\sigma^{j}$ denotes the normalized deviation of compared outputs: $\sigma^{j}=\sqrt{\frac{\sum_{y\in\{y_{i}\mid(y_{1,i},y_{2,i})\in\mathcal{P}_{m},y_{i}\in\{y_{1,i},y_{2,i}\}\}}(g^{j}(y)-\mu^{j})^{2}}{|\mathcal{P}_{m}|}}$, where $\mu^{j}$ denotes the average predicted utility of model $g^{j}$ over all outputs in $\mathcal{P}_{m}$.
@@
Figure 2 shows two surrogate models trained with the same data: a GP and our MoNNE (detailed descriptions of these models are provided in Section 5). The left-hand side of Figure 2 displays two distinct 2D utility functions. The first utility (top) is a linear utility function defined as $g_{\text{true}}(y_{1},y_{2})=y_{1}+1.5y_{2}$ and the second (bottom) is more complicated, defined as $g_{\text{true}}(y_{1},y_{2})=\sqrt{y_{1}}+0.9\text{sin}(y_{1})+\ln(y_{2}+e^{y_{1}})+4.5\sqrt{y_{2}}$. Both functions increase monotonically in the range $0,10]^{2}$. For clearer visualization, we present slices of these functions. The $x$-axis represents either $y_{1}$ (value of the first dimension) or $y_{2}$ (value of the second dimension), as indicated in the legend. For the $y_{1}+y_{2}=10$’ slice, the $x$-axis shows $y_{1}$ values while $y_{2}=10-y_{1}$. For $y_{2}=2.5$’, the $x$-axis represents $y_{1}$ values with $y_{2}$ fixed at $2.5$. Similarly, for $y_{1}=2.5$’, the $x$-axis shows $y_{2}$ values with $y_{1}$ fixed at $2.5$. Using 15 randomly chosen pairs from $y_{1}+y_{2}=10$, we constructed the GP and MoNNE models. The middle plot illustrates the GP surrogate model, while the right plot depicts the MoNNE surrogate model. As evident in Figure 2, the MoNNE model more accurately captures the underlying utility function with only 15 output pairs. Furthermore, by incorporating monotonicity constraints, MoNNE maintains a strictly increasing relationship for legends $y_{1}=2.5$’ and $y_{2}=2.5$’, whereas the GP model exhibits non-monotonic behavior.
@@ 4.2.1 Experimentation
After building the two surrogate models $f$ and $g$, we adopt qNEIUU (Astudillo and Frazier, 2020; Lin et al., 2022) to choose the next solution to be evaluated for the Experimentation stage. Following Lin et al. (2022), we calculate the qNEIUU by Monte Carlo simulation:
@@ 4.2.2 Preference Exploration
Our experiments demonstrate that IEUBO generally outperforms EUBO in the presence of utility noise. Detailed experimental results are provided in Appendix A.8. This performance difference may be attributed to how noise affects DM preferences: under noisy conditions, the base model of MoNNE might learn incorrect correlations between outputs, potentially making the independence assumption of IEUBO more robust than EUBO’s consideration of correlations. Hence, we use IEUBO as the acquisition function.
@@ 5 Experimental Results
In this section, we compare the proposed method with benchmark algorithms on six test problems. The performance of different algorithms is measured by simple regret, defined as the gap between the true optimal utility and the best utility over the observations: $\text{max}_{x\in\mathcal{X}}\{g_{\text{true}}\left(f_{\text{true}}(x)\right)\}-\text{max}_{x\in{\{x_{i}|(x_{i},y_{i})\in\mathcal{D}_{n}\}}}\{g_{\text{true}}\left(f_{\text{true}}(x)\right)\}$.
We use all four multi-objective problems that were used as benchmarks in (Lin et al., 2022), plus two additional ones. We test a wide range of utility functions from Lin et al. (2022); Astudillo and Frazier (2020), plus the Cobb-Douglas function, which is a standard function in economics, and LinearExponential, a sum of a linear component and an exponential product. Details can be found in Appendix A.5. Following the experiment setting of Lin et al. (2022), the initial number of output observations, that is, the initial dataset, is 16 (for $d<5$) or 32 (for $d\geq 5$), and the number of initial pairwise comparisons is 10 (for $d<5$) or 20 (for $d\geq 5$). In each iteration, we evaluate one solution, observing $f_{\text{true}}(x)$ and ask the DM for his preference over one pair of outputs. We compare the proposed method against the following benchmark algorithms: Random, qNParEGO (Daulton et al., 2020), qNEHVI (Daulton et al., 2020), PBO (Astudillo et al., 2023), BOPE-Linear (using a linear model as the utility surrogate model), BOPE-GP (Lin et al., 2022), BOPE-MoNNE (Ours), BOPE-BMNN (Ours, using Variational Inference-trained Bayesian Monotonic Neural Network), Known Utility (the utility function is known, so the problem becomes a Composite BO (Astudillo and Frazier, 2019)). For more details about benchmark algorithms, see Appendix A.5.
We allow for noise in DM’s decision, as discussed in Section 3. By default, $\sigma_{\text{noise}}$ is set to $0.1$ (noise is a setting of the test problem, not the algorithm). Although $0.1$ may seem small, considering that BOPE-MoNNE can achieve regret less than $0.001$ as shown in Figure 3, $0.1$ utility noise can cause enough trouble (Figure 21 in Appendix A.9 shows how many preference errors $0.1$ utility noise can make). We also conduct tests with $\sigma_{\text{noise}}=0$ and $\sigma_{\text{noise}}=0.5$ to evaluate how this noise affects the performance of MoNNEs and GP surrogate models, and the result is shown in Figure 22 in Appendix A.9.
@@
Figure 3 compares the performance of different methods, with subplots titled ‘A+B’ (A: output function, B: utility function). Excluding Known Utility algorithm, among all the BOPE algorithms, BOPE-MoNNE consistently achieves the best performance, followed by BOPE-BMNN and BOPE-GP, aligning with the uncertainty quality of these models reported in Table 1 (Appendix A.5). We evaluated uncertainty quality using a modified cross-entropy loss function: $\mathbbm{1}_{p=1}\cdot\mathbb{P}(p=1)+\mathbbm{1}_{p=-1}\cdot\mathbb{P}(p=-1)$, where $\mathbb{P}(p=1)$ and $\mathbb{P}(p=-1)$ represent the model’s predicted probabilities for the DM’s preferences. This metric takes into account both the accuracy of the model and its uncertainty, and the larger the value, the better. From Table 1, MoNNE exhibits the highest uncertainty quality, with BNN following and GP ranking last. It should be noticed that the poor uncertainty quality of GPs is mainly due to scale invariance, while the GP prediction accuracy is still good, although lower than MoNNE and BNN (see Table 2 in Appendix A.5). The experiment details about these tables are reported in Appendix A.5. In addition, BOPE-Linear works slightly better than BOPE-MoNNE when the utility function is linear, which is not surprising. However, in the other problems, the MoNNE model consistently outperforms the linear model.
Furthermore, we find that constraining output comparisons to observed values diminishes BOPE-GP’s performance (see Figure 19 in Appendix A.8). While removing this constraint improves BOPE-GP’s effectiveness, BOPE-GP still underperforms compared to BOPE-MoNNE.